# Membership Inference Attacks on LLM-based Recommender Systems

### v1

@@ Abstract
Recommender systems (RecSys) have become an essential component of many web applications. The core of the system is a recommendation model trained on highly sensitive user-item interaction data. While privacy-enhancing techniques are actively studied in the research community, the real-world model development still depends on minimal privacy protection, e.g., via controlled access. Users of such systems should have the right to choose not to share highly sensitive interactions. However, there is no method allowing the user to know which interactions are more sensitive than others. Thus, quantifying the privacy risk of RecSys training data is a critical step to enabling privacy-aware RecSys model development and deployment. We propose a membership-inference attack (MIA)- based privacy scoring method, RecPS, to measure privacy risks at both the interaction and user levels. The RecPS interaction-level score definition is motivated and derived from differential privacy, which is then extended to the user-level scoring method. A critical component is the interaction-level MIA method RecLiRA, which gives high-quality membership estimation. We have conducted extensive experiments on well-known benchmark datasets and RecSys models to show the unique features and benefits of RecPS scoring in risk assessment and RecSys model unlearning.
1 Introduction
@@
While ICL offers substantial advantages through the power of prompts, its integration into language models raises a critical issue: privacy leakage via prompts [39]. It is well known that the success of recommendation systems is based on user-interaction data, which often contains private information about user preferences, activities, and social contexts [9, 42, 46, 38]. In few-shot ICL RecSys, such data is used by RecSys to design personalized system prompts. Figure 1 shows a typical example, where the user’s sensitive historical interactions marked in red.
@@
The remaining sections include a discussion on the related work focusing on MIAs on RecSys and LLMs (Section 2), the preliminaries about the proposed methods describing how ICL-LLM-powered RecSys works (Section 3), the new threat model (Section 4), the four MIA methods (Section 5), the experimental evaluation exploring the attack performance and factors affecting these attacks (Section 6), and a conclusion about this study (Section 7).
2 Related Work
@@ 2.2 MIA on RecSys
3 Preliminaries
In this section, we introduce In-context learning RecSys in Section 3.1
3.1 In-context learning(ICL) RecSys
@@
4 THREAT MODEL
4.1 Adversary’s Objective
@@ 4.2 Adversary’s Capabilities
5 ATTACK METHODS
5.1 Direct Inquiry Attack
@@ 6.1 Experiment setup
Table 1: Statistics of datasets.
Datasets. We assess the impact of our attacks on MovieLens-1M [21] and Amazon Digital Music [26]. The statistic of the dataset is shown in Table 1. For hallucination attack, we use IMDB [1] on the MovieLens-1M as the corpus to decide whether LLM’s recommended item is hallucinated item, which aligned with the previous research [25]. Also, we find the MusicBrainz [2] can be used as corpus on the Amazon Digital Music to decide whether the LLM’s output is hallucinated item. The recommendation-specific prompts are designed using the template provided by Liu et al. [30] known for its verified performance. It is worth noting that our objective is to determine if the target user information is used in the system prompt. However, it is likely that the LLM has seen relevant text samples in the original LLM pre-training data, which we cannot distinguish. However, as demonstrated in this study [39], LLMs exhibit significantly weaker memorization about pre-training data, compared to the data in prompts.

### v2

@@ Abstract
Large language models (LLMs) based Recommender Systems (RecSys) can flexibly adapt recommendation systems to different domains. It utilizes in-context learning (ICL), i.e., the prompts, to customize the recommendation functions, which include sensitive historical user-specific item interactions, e.g., implicit feedback like clicked items or explicit product reviews. Such private information may be exposed to novel privacy attack. However, no study has been done on this important issue. We design four membership inference attacks (MIAs), aiming to reveal whether victims’ historical interactions have been used by system prompts. They are direct inquiry, hallucination, similarity, and poisoning attacks, each of which utilizes the unique features of LLMs or RecSys. We have carefully evaluated them on three LLMs that have been used to develop ICL-LLM RecSys and two well-known RecSys benchmark datasets. The results confirm that the MIA threat on LLM RecSys is realistic: direct inquiry and poisoning attacks showing significantly high attack advantages. We have also analyzed the factors affecting these attacks, such as the number of shots in system prompts and the position of the victim in the shots.
I Introduction
@@
While ICL offers substantial advantages through the power of prompts, its integration into language models raises a critical issue: privacy leakage via prompts [40]. It is well known that the success of recommendation systems is based on user-interaction data, which often contains private information about user preferences, activities, and social contexts [9, 43, 47, 39, 24, 20]. In few-shot ICL RecSys, such data is used by RecSys to design personalized system prompts. Figure 1 shows a typical example, where the user’s sensitive historical interactions marked in red.
@@
The remaining sections include a discussion on the related work focusing on MIAs on RecSys and LLMs (Section II), the preliminaries about the proposed methods describing how ICL-LLM-powered RecSys works (Section III), the new threat model (Section IV), the four MIA methods (Section V), the experimental evaluation exploring the attack performance and factors affecting these attacks (Section VI), and a conclusion about this study (Section VII).
II Related Work
@@ II-B MIA on RecSys
III Preliminaries
In this section, we introduce In-context learning RecSys in Section III-A
III-A In-context learning(ICL) RecSys
@@
IV THREAT MODEL
IV-A Adversary’s Objective
@@ IV-B Adversary’s Capabilities
V ATTACK METHODS
V-A Direct Inquiry Attack
@@ VI-A Experiment setup
TABLE I: Statistics of datasets.
Datasets. We assess the impact of our attacks on MovieLens-1M [22] and Amazon Digital Music [27]. The statistic of the dataset is shown in Table I. For hallucination attack, we use IMDB [1] on the MovieLens-1M as the corpus to decide whether LLM’s recommended item is hallucinated item, which aligned with the previous research [26]. Also, we find the MusicBrainz [2] can be used as corpus on the Amazon Digital Music to decide whether the LLM’s output is hallucinated item. The recommendation-specific prompts are designed using the template provided by Liu et al. [31] known for its verified performance. It is worth noting that our objective is to determine if the target user information is used in the system prompt. However, it is likely that the LLM has seen relevant text samples in the original LLM pre-training data, which we cannot distinguish. However, as demonstrated in this study [40], LLMs exhibit significantly weaker memorization about pre-training data, compared to the data in prompts.